# Towards Universal Performance Modeling for Machine Learning Training on Multi-GPU Platforms

### v2

@@ Abstract
Characterizing and predicting the training performance of modern machine learning (ML) workloads on compute systems with compute and communication spread between CPUs, GPUs, and network devices is not only the key to optimization and planning but also a complex goal to achieve. The primary challenges include the complexity of synchronization and load balancing between CPUs and GPUs, the variance in input data distribution, and the use of different communication devices and topologies (e.g., NVLink, PCIe, network cards) that connect multiple compute devices, coupled with the desire for flexible training configurations. Built on top of our prior work for single-GPU platforms, we address these challenges and enable multi-GPU performance modeling by incorporating (1) data-distribution-aware performance models for embedding table lookup, and (2) data movement prediction of communication collectives, into our upgraded performance modeling pipeline equipped with inter-and intra-rank synchronization for ML workloads trained on multi-GPU platforms. Beyond accurately predicting the per-iteration training time of DLRM models with random configurations with a geomean error of 5.21% on two multi-GPU platforms, our prediction pipeline generalizes well to other types of ML workloads, such as Transformer-based NLP models with a geomean error of 3.00%. Moreover, even without actually running ML workloads like DLRMs on the hardware, it is capable of generating insights such as quickly selecting the fastest embedding table sharding configuration (with a success rate of 85%).
@@ I Introduction
Modern machine learning (ML) workloads tend to grow in both size and computation, usually beyond the capability of one single GPU to host and train. Platforms with multiple compute devices such as GPUs, therefore, play a essential role in undertaking training jobs for these workloads. Modern industrial recommendation models, such as DLRM [1], consist of embedding tables of hundreds of gigabytes and are trained in a distributed way on hierarchical compute systems, e.g., 16 nodes $\times$ 8 GPUs, as an advanced software-hardware co-designed system [2]. In the area of large language models (LLM), the rapidly growing number of model parameters requires an equally rapid increase in the number of compute devices used to train them. Examples include GPT3 [3] (2020, 175B (parameter count, same below), 1024 NVIDIA V100 GPUs (estimated)) Megatron-Turing NLG [4] (2022, 530B, 2240 80-GB NVIDIA A100 GPUs), PaLM [5] (2022, 540B, 6144 Google TPU v4), OPT-175B [6] (2022, 175B, 992 80-GB NVIDIA A100 GPUs), and LLaMA [7] (2023, 65B, 2048 80-GB NVIDIA A100 GPUs). For such daunting multi-GPU training jobs, ML practitioners are interested in not only training performance (e.g., prediction of iteration time, query-per-sec (QPS), FLOPS, etc.) but also how to speed up training and use the hardware most efficiently. Characterizing the performance behavior of such multi-GPU training jobs is the key to identifying and optimizing performance bottlenecks, and will subsequently help ML practitioners save development time and budget, improve ML model quality provided to users, and benefit the environment by avoiding excessive emissions of CO2.
@@
  • More importantly, the synchronization behaviors of multiple GPU streams on the same device or across multiple devices are complicated. Previous works have only inadequate or premature modeling for these behaviors, so they are not sufficient to account for all the GPU idle time on each rank caused by data dependencies and multi-GPU synchronization, and as a result, cannot accurately model workload execution time.
@@
Also, even in single-GPU setups, the limitations of the performance modeling of certain operators (ops) in prior work inhibit us from extending it to a broader range of workloads. As an important example, embedding lookups in real-world DLRM pipelines have unknown input distributions and loads across batches and devices, resulting in unpredictable memory access patterns and thus difficulty in predicting their performance. Existing embedding lookup performance models are rigid and only able to cover limited problem sizes. Besides, extending the performance model to other workloads such as natural language processing (NLP) models requires the support of extra element-wise or minor ops.
@@
To summarize, the performance modeling problems of communication collectives, multi-GPU stream synchronization, embedding lookups with randomly distributed input data, and additional minor ops such as layer_norm and dropout remained unsolved until we address them in this paper. We extend our previous performance-modeling work [14] and inherit its critical-path-based execution trace simulation method to model the training performance of deep learning recommendation models (DLRM) and natural language processing (NLP) workloads on multi-GPU platforms. The novel and major contributions include:
@@
  • (Section III-C) Added performance models for communication operations (all-to-all and all-reduce) by improving a basic heuristic model with straightforward and efficient sigmoid curve fitting.
@@
  • (Section III-D) An enhanced critical-path-based end-to-end (E2E) performance modeling algorithm.
@@
  • (Section III-E) Improved embedding-table-lookup performance modeling with flexible lookup numbers and patterns (i.e., input data distribution) through an ML-based approach.
@@
  • (Section III-F) Supported extra minor ops such as layer_norm and dropout for NLP models.
@@
We claim that both inter- and intra-rank synchronizations are the keys to accurately modeling ML workloads training performance on multi-GPU, or even more broadly, all types of workloads running on multi-heterogeneous-device platforms. Therefore, the second item above is the most critical contribution of this paper. Our performance model obtains 5.21% and 3.00% geomean prediction errors on predicting the per-iteration performance of randomly generated industrial-scale DLRM workloads and Transformer-based NLP models based on BERT [15], GPT2 [16], and XLNet [17], respectively, on two different multi-GPU platforms. In addition, in a use case of choosing a (DLRM) embedding table sharding configuration without running the workloads, our performance model achieves an 85% success rate in selecting either the fastest config or predicting the execution time with less than 10% error, showcasing our performance model’s ability to generate insights for multi-GPU training optimization.
@@ II-A Performance Modeling for Both Single-Device and Distributed Training
There is abundant prior art in predicting the execution time of per-batch training and/or inference of ML workloads on a single GPU. A majority of them [8, 9, 10, 11, 12, 13, 14] focus on convolutional neural networks (CNNs), while some also cover NLP [12] and recommendation models like DLRM [14]. These works developed methodologies for accurately modeling performance on the op/kernel levels, though in general did not address the challenges of multi-GPU or distributed training.
@@
Early performance models for distributed training are mostly analytical and consider computation and communication separately. Yan et al. [18] studied data and model parallelism of CNNs on CPU clusters and modeled their performance and scalability. Oyama et al. [19] proposed a performance model for GPU-equipped supercomputers, which predicts per-batch execution time with low error based on statistics of mini-batch size and staleness for an asynchronous SGD algorithm. It also provides some generality in terms of precision and communication standards. As a use case, both approaches claimed to be able to search for the best system configurations for distributed training of CNNs. Qi et al. [20] introduced PALEO for per-layer CNN execution time estimation, and demonstrated its ability to accurately model CNN training performance at scale on cloud clusters.
@@
However, these prior successes do not extend well to modern ML workloads. Today, the implementation of ML workloads as well as the environment and setup of training clusters have evolved significantly and thus outdated features such as sub-batch size (fewer than 16 [19] vs. considerably larger sizes in modern settings, e.g., 2048 to 65536), model workload size and diversity, and hardware settings. Prior work either has a lack of support on GPUs [18], or a complex methodology that inhibits generalizing to other pipelines [19]. As decentralized distributed training with multiple GPUs is increasingly preferred to parameter servers, the problem of modeling the communication among GPUs and subsequently accounting for interaction and overlap between GPU streams must also be addressed for accurate prediction. This rules out per-layer methods such as Qi et al. [20].
@@
Wang et al. [21] implied a possible speed-of-light (SOL) model that the total execution time is the maximum among the time of data, computation, and communication, assuming these operations are perfectly overlapped. Inspired by this, and thanks to the detailed information we obtain from the execution trace [22], we further predict the training performance in general when these operations are not perfectly overlapped. Yang et al. [23] proposed the computational-graph simulation-based PerfEstimator running on one single device for distributed training of CV models, with computation-communication overlap estimated by an analytical scaling factor. In our methods, this overlap is estimated by critical-path-based methods that is less error-prone especially when workloads across devices are imbalanced.
@@ II-B Model Parallelism and Sharding
As modern ML workloads have grown rapidly in size, exploiting model parallelism has become a critical tool in scaling up ML workloads. Lepikhin et al. [24] summarize the main challenges of model parallelism, which include: 1) device underutilization due to the sequential dependency of the network and gradient-based optimization; 2) superlinear compute cost vs. model size; 3) poor infra scalability; 4) non-trivial implementation of partitioning strategies. As an example of one increasingly common use of model parallelism, modern recommendation models such as industrial DLRMs often have hundreds of embedding lookup tables with up to tens of millions of rows, which are too large to be stored on a single GPU and thus must be sharded and dispatched to multiple GPUs. Load-balancing these sharded tables (to achieve uniform access latency and memory usage) thus becomes a critical problem, one which has been addressed by previous researchers. Lui et al. [25] first applied baseline (e.g., dimension-based, row-based, and size-based) heuristics to shard embedding tables in DLRMs. Sethi et al. proposed RecShard [26] that uses integer linear programming (ILP) to optimize the sharding problem. Zha et al. further improved the sharding efficiency and balance by addressing the problem with reinforcement learning in AutoShard [27] and Dreamshard [28]. In our paper, we demonstrate in a case study that our performance model is able to evaluate the performance of multiple sharding algorithms and quickly select the one that leads to the best E2E execution time for DLRM training. Notice that we are not proposing a smart sharding algorithm here as these previous works did.
@@ III Methodology
Figure 1: An overview of our prediction pipeline, based on the one proposed by Lin et al. [14]. We mark new components with small darker shapes and italic text, such as microbenchmark and kernel-level performance models of FBGEMM embedding lookup (FEL), all-to-all (A2A), and all-reduce (AR), as well as inter- and intra-rank synchronization mechanism in the critical-path algorithm to handle multi-GPU end-to-end performance prediction.
Our previous work on single-GPU performance modeling [14] is both modular and extendable. We leverage and extend it with a focus on the performance modeling on single-node multi-GPU platforms. The reason why this makes sense is because kernel performance models are reused and shared across different ML workloads as inputs, while the E2E time prediction by simulating the execution trace of the target workload and traversing ops in it is dictated by the same critical-path-based algorithm that is both model-architecture-agnostic and platform-agnostic. In this current work, we extend this E2E performance model for new ML workloads and multi-GPU platforms by both adding kernel performance models for dominating ops/kernels of these new workloads and addressing multi-GPU execution in the E2E simulation.
Figure 1 depicts an overview of our prediction pipeline. It inherits its main modules and procedures, and reuses both all kernel performance models (for GEMM, memcpy, transpose, etc) and the overhead estimator from the original pipeline. The logic of the prediction pipeline remains unchanged. To accumulate system assets (blue cylinders in Figure 1), we profile ML workloads written with PyTorch, extract execution traces, and analyze traces for dominating ops. Then, we collect microbenchmark data for target kernels/ops, design/train and verify performance models using this data, and extract overhead statistics for these kernels/ops as well. In the prediction phase, the pipeline takes the execution traces of a workload as the input, simulates the execution by traversing the ops in the trace, and predicts the per-iteration training time of the workload with the system assets within several seconds. We focus on DLRM and Transformer-based NLP workloads in this work, but we expect our techniques can seamlessly extend and apply to other ML workloads (details discussed in Section V-B).
@@ III-A Dataset Exploration
The open-source DLRM dataset [29] contains synthetic embedding lookup data that resembles the memory-access reuse pattern of Meta’s production data. The 2021 data includes a batch of 65536 samples embedding lookup data of 856 tables, while the 2022 data includes a batch of 131072 samples of 788 tables. Due to the unpredictable lookup pattern in production data, the pooling factors (average number of lookups per data sample, denoted with $L$) vary significantly across all tables. This is very different from what has been studied in previous work [30, 14], which considers $L$ as a fixed value across data batches and even tables in the same workload. Our exploratory analysis (Figure 3) shows that average $L$ values of tables in the dataset tend to concentrate in the range of $[0,10)$, especially $[0,1]$, while there are still a number of tables with average $L$ value spreading to a few hundred. We define tables with an average $L$ equal to greater than 20 as “heavy tables”. As shown in Figure 3, there are 103 and 166 heavy tables in the 2021 and 2022 data respectively.
@@
Figure 3: Histogram of average $L$ values of embedding lookup tables in the dataset. All bins but the last one are right-open.
@@ III-B Multi-GPU Training Benchmark and Analysis
We first benchmark several ML workloads that will be further analyzed in later sections to understand the behavior of multi-GPU training performance. Figure 2 shows the per-GPU-stream execution time breakdown of three workloads: DLRM_Heavy_EL with a batch size of 4096 (Task_14 in Section IV-B1, involving heavy embedding tables only), DLRM_Light_EL with a batch size of 4096 (Task_5 in Section IV-B1, involving random embedding tables), and GPT2 with a batch size of 64 on each rank of the 4xGV100 platform (details shown in Section IV). The runtime is contributed by a few groups of kernels and types: embedding lookup (forward and backward), GEMM (forward and backward), memory operations (concat, copy, transpose, etc), others (normalizations, elementwise, etc), communication (all-to-all and all-reduce), and idle time. We summarize the following insights from our benchmarking:
@@
  • Communication operations can be but are not always a performance bottleneck. It is obvious that communication operations dominate the two DLRM models. This is mainly because in addition to the normal all-reduce caused by accumulating gradients in the backward pass, training the embedding lookup module in DLRM with model-parallelism introduces the all-to-all operations for merging data with that from the bottom and top multilayer perceptron (MLP) modules trained with data parallelism. In most of the stream pairs shown here, except for DLRM_Heavy_EL’s rank 2, the communication stream has a longer GPU active time than that of the computation stream, which means it can never be fully overlapped by the computation and thus similar workloads tend to thrash the communication network of the interconnected multi-GPU platform. In contrast, GPT2 is quite compute-dominated; all-reduce time is minor and can theoretically be fully overlapped.
@@
  • Load balance substantially varies across the workloads. DLRM models experience load imbalance across ranks due to the existence of embedding lookup tables with various compute and communication loads. Notice that instead of the forward and backward computation of embedding lookup itself, it is in fact the all-to-all it introduces that dominates, especially when the embedding lookup is light. Examples such as DLRM_Light_EL’s rank 3 and 4 show that communication can occupy up to 80% of the execution time, in which case GPU computation resources are wasted by sitting idle. In contrast, we see an almost-perfect load balance for GPT2 as it is purely DDP-trained and the load is evenly distributed.
@@
  • The significance of GPU stream idle time is way more than “being non-negligible”. What we see in single-GPU training is that idle time is contributed mostly by CPU op calls and overheads that block the scheduling of GPU kernels; data dependencies play a less important role given there is only one stream. On the contrary, in the case of multi-GPU training data dependencies of ops lead to streams that wait on each other and thus dominate waiting times. This time cannot be predicted in the same way (i.e., statistically) as the previous work such as Lin et al. [14] did. As we show later, the kernel execution time on each stream varies and thus the overlap of these kernels on different GPU streams is complicated and case-by-case. Therefore, to understand and simulate how GPU streams interact and synchronize with each other is the key to accurately predicting the per-iteration time of training.
@@
  • As we support more kinds of ML workloads such as (Transformer-based) NLP models, we will see an increasing number of ops that must be modeled accurately to get an accurate result. We see that “other” ops (layer-norm, dropout, gelu, tanh, etc.) contribute to more than 20% of execution time on GPT2, although each one of them only contributes a tiny amount of time every time it is executed. This might not be a surprise, given the performance of Transformer-based models has been profoundly studied these years. Techniques such as advanced layer fusion may be able to effectively optimize and reduce their execution time, but that is beyond the scope of this paper. Under our experiment settings and from the performance-modeling point of view, this observation basically means more ops from this category should be supported than the previous work did.
@@
The lessons from the above two subsections guide the direction of our research, which includes the following novel components or improvements of our performance model:
@@
  • Advanced E2E performance modeling for multi-GPU training.
@@
III-C Communication Collective Performance Modeling

### v3

@@ Abstract
Characterizing and predicting the training performance of modern machine learning (ML) workloads on compute systems with compute and communication spread between CPUs, GPUs, and network devices is not only the key to optimization and planning but also a complex goal to achieve. The primary challenges include the complexity of synchronization and load balancing between CPUs and GPUs, the variance in input data distribution, and the use of different communication devices and topologies (e.g., NVLink, PCIe, network cards) that connect multiple compute devices, coupled with the desire for flexible training configurations. Built on top of our prior work for single-GPU platforms, we address these challenges and enable multi-GPU performance modeling [^1] Code is open-sourced at https://github.com/owensgroup/ml_perf_model. by incorporating (1) data-distribution-aware performance models for embedding table lookup, and (2) data movement prediction of communication collectives, into our upgraded performance modeling pipeline equipped with inter-and intra-rank synchronization for ML workloads trained on multi-GPU platforms. Beyond accurately predicting the per-iteration training time of deep learning recommendation models (DLRM) models with random configurations with a geomean error of 5.21% on two multi-GPU platforms, our prediction pipeline generalizes well to other types of ML workloads, such as Transformer-based natural language processing (NLP) models with a geomean error of 3.00%. Moreover, even without actually running ML workloads like DLRMs on the hardware, it is capable of generating insights such as quickly selecting the fastest embedding table sharding configuration (with a success rate of 85%).
@@ I Introduction
Modern machine learning (ML) workloads tend to grow in size and computation, usually beyond the capability of one single GPU to host and train. Therefore, platforms with multiple compute devices such as GPUs are critical in undertaking training jobs for these workloads. Modern industrial recommendation models, such as DLRM [1], consist of embedding tables of hundreds of gigabytes and are trained in a distributed way on hierarchical compute systems, e.g., 16 nodes $\times$ 8 GPUs, as an advanced software-hardware co-designed system [2]. For large language models (LLM), the rapidly growing number of model parameters requires an equally rapid increase in the number of compute devices used to train them. Examples include GPT3 [3] (2020, 175B (parameter count, same below), 1024 NVIDIA V100 GPUs (estimated)) Megatron-Turing NLG [4] (2022, 530B, 2240 80-GB NVIDIA A100 GPUs), PaLM [5] (2022, 540B, 6144 Google TPU v4), OPT-175B [6] (2022, 175B, 992 80-GB NVIDIA A100 GPUs), and LLaMA [7] (2023, 65B, 2048 80-GB NVIDIA A100 GPUs). For such daunting multi-GPU training jobs, ML practitioners are interested in not only the training performance (e.g., iteration time, query-per-sec (QPS), FLOPS, etc.) but also how to speed up training and use the hardware most efficiently. Characterizing the performance behavior of such multi-GPU jobs is the key to identifying and optimizing performance bottlenecks, and subsequently helping ML practitioners save development time and budget, improve ML model quality provided to users, and benefit the environment by avoiding excessive emissions of CO2.
@@
  • More importantly, the synchronization behaviors of multiple GPU streams on the same device or across multiple devices are complicated. Previous works have only inadequate or premature modeling for these behaviors, making them insufficient to account for all the GPU idle time on each rank caused by data dependencies and multi-GPU synchronization, and as a result, cannot accurately model workload execution time.
@@
Also, even in single-GPU setups, the limitations of the performance modeling of certain operators (ops) in prior work inhibit us from extending it to a broader range of workloads. As an important example, embedding lookups in real-world DLRM pipelines have unknown input distributions and loads across batches and devices, resulting in unpredictable memory access patterns and thus difficulty in predicting their performance. Existing embedding lookup performance models are rigid and can only cover limited problem sizes. Besides, extending the performance model to other workloads such as natural language processing (NLP) models requires the support of extra element-wise or minor ops.
@@
To summarize, the performance modeling problems of communication collectives, multi-GPU stream synchronization, embedding lookups with randomly distributed input data, and additional minor ops such as layer_norm and dropout remained unsolved until we address them in this paper. We extend our previous performance-modeling work [14] and inherit its critical-path-based execution trace simulation method to model the training performance of DLRM and NLP workloads on multi-GPU platforms. The novel and major contributions include:
@@
  • (Section IV-A) Added performance models for communication operations (all-to-all and all-reduce) by improving a basic heuristic model with straightforward and efficient sigmoid curve fitting.
@@
  • (Section IV-B) An enhanced critical-path-based end-to-end (E2E) performance modeling algorithm.
@@
  • (Section IV-C) Improved embedding-table-lookup performance modeling with flexible lookup numbers and patterns (i.e., input data distribution) through an ML-based approach.
@@
  • (Section IV-D) Supported extra minor ops such as layer_norm and dropout for NLP models.
@@
We claim that the modeling of both inter- and intra-rank synchronizations are the keys to accurately modeling ML workloads training performance on multi-GPU, or even more broadly, all types of workloads running on multi-heterogeneous-device platforms. Therefore, the proposal of the critical-path-based end-to-end (E2E) performance modeling algorithm (second item above) is the most significant contribution of this paper. Our performance model obtains 5.21% and 3.00% geomean prediction errors on predicting the per-iteration performance of randomly generated industrial-scale DLRM workloads and Transformer-based NLP models based on BERT [15], GPT2 [16], and XLNet [17], respectively, on two different multi-GPU platforms. In addition, in a use case of choosing a (DLRM) embedding table sharding configuration without running the workloads, our performance model achieves an 85% success rate in selecting either the fastest config or predicting the execution time with less than 10% error, showcasing our performance model’s ability to generate insights for multi-GPU training optimization.
@@ II-A Performance Modeling for Both Single-Device and Distributed Training
There is abundant prior art in predicting the execution time of per-batch training and/or inference of ML workloads on a single GPU. A majority of them [8, 9, 10, 11, 12, 13, 14] focus on convolutional neural networks (CNNs), while some also cover NLP [12] and recommendation models like DLRM [14]. These works developed methodologies for accurately modeling performance on the op/kernel levels, though they did not address the challenges of multi-GPU or distributed training in general.
@@
Early performance models for distributed training are mostly analytical and consider computation and communication separately. Yan et al. [18] studied data and model parallelism of CNNs on CPU clusters and modeled their performance and scalability. Oyama et al. [19] predicted per-batch execution time for GPU-equipped supercomputers with low error based on statistics of mini-batch size and staleness for an asynchronous SGD algorithm. It also provided some generality in terms of precision and communication standards. As a use case, both approaches claimed to be able to search for the best system configurations for distributed training of CNNs. Qi et al. [20] introduced PALEO for per-layer CNN execution time estimation and demonstrated its ability to accurately model CNN training performance at scale on cloud clusters.
@@
However, these prior successes do not extend well to modern ML workloads. Today, the implementation of ML workloads as well as the environment and setup of training clusters have evolved significantly. This outdates features such as sub-batch size (fewer than 16 [19] vs. considerably larger sizes in modern settings, e.g., 2048 to 65536), model workload size and diversity, and hardware settings. Prior work either lacks support on GPUs [18] or has a complex methodology that inhibits generalizing to other pipelines [19]. As decentralized distributed training such as DDP (distributed data-parallel) with multiple GPUs is increasingly preferred to parameter servers, modeling the communication among GPUs to account for interaction and overlap between GPU streams must also be addressed for accurate prediction. This rules out per-layer methods such as Qi et al. [20].
@@
Wang et al. [21] implied a possible speed-of-light (SOL) model that the total execution time is the maximum among the time of data, computation, and communication, assuming these operations are perfectly overlapped. Inspired by this, and thanks to the detailed information we obtain from the execution trace [22], we further predict the training performance in general when these operations are not perfectly overlapped. Yang et al. [23] proposed the computational-graph simulation-based PerfEstimator running on a single device for distributed training of CV models, with computation-communication overlap estimated by an analytical scaling factor. In our methods, this overlap is estimated by the critical-path-based method which is less error-prone especially when workloads across devices are imbalanced.
@@ II-B Model Parallelism and Sharding
As modern ML workloads have grown rapidly in size, exploiting model parallelism has become a critical tool in scaling up ML workloads. Lepikhin et al. [24] summarize the main challenges of model parallelism, which include: 1) device underutilization due to the sequential dependency of the network and gradient-based optimization; 2) superlinear compute cost vs. model size; 3) poor infra scalability; 4) non-trivial implementation of partitioning strategies. As an example of one increasingly common use of model parallelism, modern recommendation models such as industrial DLRMs often have hundreds of embedding lookup tables with up to tens of millions of rows, which are too large to be stored on a single GPU and thus must be sharded and dispatched to multiple GPUs. Balancing the load of these sharded tables (to achieve uniform access latency and memory usage) thus becomes a critical problem that previous researchers have tried to solve in different ways. Lui et al. [25] first applied baseline (e.g., dimension-based, row-based, and size-based) heuristics to shard embedding tables in DLRMs. Sethi et al. proposed RecShard [26] that uses integer linear programming (ILP) to optimize the sharding problem. Zha et al. further improved the sharding efficiency and balance by the reinforcement-learning-based AutoShard [27] and Dreamshard [28]. In our paper, we demonstrate through a case study that our performance model can evaluate the performance of multiple sharding algorithms and quickly select the one that leads to the best E2E execution time for DLRM training. Notice that we are not proposing a smart sharding algorithm here as these previous works did.
@@ III Methodology
Figure 1: An overview of our prediction pipeline, based on the one proposed by Lin et al. [14]. We mark new components with small darker shapes and italic text, such as microbenchmark and kernel-level performance models of FBGEMM embedding lookup (Section IV-C), all-to-all (A2A), and all-reduce (AR) (Section IV-A), as well as inter- and intra-rank synchronization mechanism in the critical-path algorithm (Section IV-B) to handle multi-GPU end-to-end performance prediction.
Figure 1 provides an overview of our multi-GPU performance modeling pipeline with the modules that exist in our previous work [14] marked by light blocks and extended modules of this current work marked by dark blocks and italic text. The previous work builds a solid foundation for the pipeline. In the analysis track, we profile ML workloads written with PyTorch and analyze the profiler traces to collect dominating kernels/ops. Then we collect microbenchmark data for these kernels/ops, design/train and verify performance models using this data, and extract the overhead statistics for these kernels/ops. The kernel benchmark data, kernel performance models, and overhead statistics are saved as the assets of the pipeline ((blue cylinders in Figure 1)). In the prediction track, the pipeline extracts the execution trace (renamed from the “execution graph”) of an input ML workload, simulates it by traversing its ops with the pipeline assets, and generates the per-iteration E2E training time prediction of the input ML workload within several seconds.
We leverage the modularity of the previous work and extend it with a focus on performance modeling on single-node multi-GPU platforms. This makes sense for two reasons: 1) the pipeline assets are reused and shared across different input ML workloads, and 2) the E2E time prediction by simulating the execution trace of the target workload and traversing ops in it is dictated by the same critical-path-based logic that is both model-architecture-agnostic and platform-agnostic. In our current work, we add kernel performance models for dominating kernels/ops of ML workloads running on multi-GPU platforms and address multi-GPU execution in the E2E simulation. The updated pipeline inherits the main modules and procedures and reuses all kernel performance models (for GEMM, memcpy, transpose, etc) and the overhead estimator from the previous version, while the critical-path-based simulation algorithm is enriched to adapt to multi-GPU execution. We focus on DLRM and Transformer-based NLP workloads in this work, but we expect our techniques can seamlessly extend and apply to other ML workloads (details discussed in Section VI-B).
@@ III-A Dataset Exploration
The open-source DLRM dataset [29] contains synthetic embedding lookup (EL) data that resembles the memory-access reuse pattern of Meta’s production data. The 2021 data includes one batch of 65536 EL data samples from 856 tables, while the 2022 data includes one batch of 131072 samples from 788 tables. Due to the unpredictable lookup pattern in production data, the pooling factors (average number of lookups per data sample, denoted with $L$) vary significantly across all tables. This is very different from the previous work [30, 14] which considers $L$ as a fixed value across data batches and even tables in the same workload. Our exploratory analysis (Figure 3) shows that average $L$ values of tables in the dataset tend to concentrate in the range of $[0,10)$, especially $[0,1]$, while there are still some tables with average $L$ value spreading to a few hundred. We define tables with an average $L$ equal to greater than 20 as “heavy tables”. As shown in Figure 3, there are 103 and 166 heavy tables in the 2021 and 2022 data respectively.
@@
Figure 3: Histogram of average $L$ values of EL tables in the dataset. All bins but the last one are right-open.
@@ III-B Multi-GPU Training Benchmark and Analysis
We first benchmark several ML workloads that will be further analyzed in later sections to understand the behavior of multi-GPU training performance. Figure 2 shows the per-GPU-stream execution time breakdown of three workloads: DLRM_Heavy_EL with a batch size of 4096 (Task_14 in Section V-B1, involving heavy embedding tables only), DLRM_Light_EL with a batch size of 4096 (Task_5 in Section V-B1, involving random embedding tables), and GPT2 with a batch size of 64 on each rank of the 4xGV100 platform (details shown in Section V). The runtime is contributed by a few groups of kernels and types: EL (forward and backward), GEMM (forward and backward), memory operations (concat, copy, transpose, etc), others (normalizations, elementwise, etc), communication (all-to-all and all-reduce), and idle time. We summarize the following insights from our benchmark:
@@
  • Communication operations could be but are not always a performance bottleneck. Obviously, communication operations dominate the two DLRM models, mainly because in addition to all-reduce commonly caused by accumulating gradients in the backward pass, training the EL module in DLRM with model-parallelism introduces all-to-all for merging data with that from the multilayer perceptron (MLP) modules trained with data parallelism. In most pairs of streams shown here (except for DLRM_Heavy_EL’s rank 2), the communication stream has a longer GPU active time than the computation stream. This means it cannot be fully overlapped by the computation, and thus similar workloads tend to thrash the communication network of the interconnected multi-GPU platform. In contrast, GPT2 is quite compute-dominated; all-reduce time is minor and can theoretically be fully overlapped.
@@
  • Load balance substantially varies across the workloads. DLRM models experience load imbalance across ranks due to the existence of EL tables with various compute and communication loads. Notice that instead of the forward and backward computation of EL itself, it is the all-to-all it introduces that dominates, especially when the EL is light. Examples such as DLRM_Light_EL’s rank 3 and 4 show that communication can occupy up to 80% of the execution time, in which case GPU computation resources are wasted by sitting idle. In contrast, we see an almost-perfect load balance for GPT2 as it is purely DDP-trained and the load is evenly distributed.
@@
  • The significance of GPU stream idle time is way more than “being non-negligible”. What we see in single-GPU training is that idle time is contributed mostly by CPU op calls and overheads that block the scheduling of GPU kernels; data dependencies play a less important role given there is only one stream. On the contrary, in the case of multi-GPU training data dependencies of ops lead to streams that wait on each other and thus dominate waiting time. This time cannot be predicted in the same way (i.e., statistically) as the previous work such as Lin et al. [14] did, because the kernel execution time on each stream varies and thus the overlap of these kernels on different GPU streams is complicated and case-by-case. Therefore, to understand and simulate how GPU streams interact and synchronize with each other is the key to accurately predicting the per-iteration time of training.
@@
  • As we support more kinds of ML workloads such as (Transformer-based) NLP models, we will see an increasing number of ops that must be modeled accurately to get an accurate result. We see that “other” ops (layer-norm, dropout, gelu, tanh, etc.) contribute to more than 20% of execution time on GPT2, although each one of them only contributes a tiny amount of time every time it is executed. This might not be a surprise, given the performance of Transformer-based models has been profoundly studied these years. Techniques such as advanced layer fusion may be able to effectively optimize and reduce their execution time, but that is beyond the scope of this paper. Under our experiment settings and from the performance-modeling point of view, this observation means more ops from this category should be supported than the previous work did.
@@
The lessons we learn guide the direction of our research, which includes the following novel components or improvements of our performance model covered in the next section:
@@
  • E2E performance modeling for multi-GPU training.
@@
IV Kernel and End-to-End Performance Modeling for ML Training on Multi-GPU
IV-A Communication Collective Performance Modeling